# Differentially Private Distribution Release of Gaussian Mixture Models via KL-Divergence Minimization

### v1

@@ Abstract
Index Terms:
Gaussian mixture model, density estimation, distribution release, model fitting, differential privacy.
@@ I Introduction
In recent years, the remarkable success of data-driven artificial intelligence (AI) has spurred an increasing demand for the sharing and analysis of large-scale, multi-class, and high-dimensional datasets across a variety of domains, such as healthcare records, consumer transactions, and mobility traces. Organizations have recognized the potential of sharing data statistics to enhance data mining, improve public services, optimize recommendations, and facilitate data simulation [bishop2006pattern]. However, sharing raw data or even their statistics raise significant privacy concerns, especially when sensitive attributes of individuals might be inferred. This tension underscores the need for privacy-preserving mechanisms that allow the release of data statistics without compromising personal or proprietary information [dwork2006calibrating].
@@
Among the various statistical modeling approaches, Gaussian Mixture Models (GMMs) stand out as a versatile tool for representing complex, multivariant, and multi-modal data distributions [figueiredo2002unsupervised]. GMMs are widely applied in data mining, machine learning (ML), and pattern recognition. By modeling the overall distribution as a mixture of several Gaussian components, GMMs naturally capture latent subgroups or clusters in the data, such as different risk profiles in healthcare or distinct spending behaviors in retail. Furthermore, GMMs provide a compact and interpretable parameterization, including component means, covariances, and mixture weights (also known as categorical frequencies), which can be shared more efficiently than raw data records. As a motivating real-world application, GMMs are accurate in representing the statistics of energy consumption data. In particular, the logarithmic values of load demand profiles in Advanced Metering Infrastructure (AMI) data have been empirically shown to fit well to GMMs [9799519]. In the context of energy usage, these load demand profiles represent the amount of electricity consumed over time by a given user or household. When fitted to a GMM, the data naturally forms clusters, with each Gaussian component capturing different usage patterns (e.g., high, medium, and low usage times).
@@
Despite their advantages, releasing GMM parameters directly still poses privacy risks, particularly when the underlying dataset contains sensitive information. An adversary could potentially exploit small mixture components or extreme values to infer specific individuals’ records or their associated class labels. For instance, an adversary could inspect the component means of the GMM model fitted to the AMI dataset and detect changes that could identify the presence of a specific customer. Small perturbations in the model, such as flipping the label of a single customer, could significantly alter the estimated parameters, making it easier for the adversary to infer private information. To mitigate these risks, differential privacy (DP) [dwork2006calibrating, dwork2014algorithmic] offers a robust framework for quantifying and controlling the privacy loss resulting from the inclusion or exclusion of any individual record in a dataset. By introducing carefully calibrated noise, we can estimate mixture model parameters that satisfy strong privacy guarantees while still enabling meaningful statistical analysis.
@@
A central challenge in applying DP mechanisms to GMM fitting is maintaining fidelity between the original non-private data distribution and the released one under a DP constraint. Conventional DP mechanisms typically add artificial noise to perturb the model parameters in order to protect data privacy. However, this perturbation inevitably compromises the accuracy of the released GMM. For instance, the noise added to the Gaussian means could distort the true cluster centers, leading to incorrect clustering and a GMM that fails to represent the underlying data structure. Existing work [7590445] addressed differentially private GMMs by adding white Laplace noise to the mixture weights, means, and covariances individually. This approach models the accuracy of differentially private GMMs using the mean-square errors (MSEs) of individual parameters.
@@
Rather than the MSEs that measure the accuracy of releasing individual parameters, we argue that distribution divergence, specifically Kullback–Leibler (KL) divergence, is another comprehensive metric for quantifying the fidelity of GMMs. KL divergence is a well-established statistical distance metric in information theory that quantifies how much one probability distribution diverges from another [cover1999elements]. When used to characterize the utility of distribution release, KL divergence measures the “information” lost when substituting the true distribution with an approximate, privacy-preserving model. It is well-known that computing maximum likelihood estimation (MLE) is asymptotically equivalent to minimizing the KL divergence between the empirical and true distributions [DL_Goodfellow]. Thus, publishing a data distribution with minimal KL divergence ensures that the released model is as accurate as possible in the maximum-likelihood sense. Moreover, when applied to data simulation and augmentation for ML, KL divergence is closely related to the generalization ability of ML models. Extensive research has established PAC-Bayes bounds for quantifying the generalization error of ML models [MAL-112], which depends critically on the KL divergence between the training and testing data distributions. In this context, KL divergence serves as a powerful metric for evaluating how well an ML model trained on a given data distribution (e.g., a released GMM under privacy constraints) generalizes to the true data, thus providing an important tool for assessing the performance of data synthesis and augmentation in ML training.
@@
KL divergence provides a unified framework for characterizing the combined impact of noise perturbation on the accuracy of the released GMM, taking into account the effects of noise on all model parameters. However, characterizing the KL divergence between the original and DP-preserved distributions, particularly when applying DP mechanisms to the released GMM parameters, remains an open problem. In this work, we show that the KL divergence for GMM parameter release exhibits a closed-form expression, thereby opening new research avenues for designing DP mechanisms that minimize KL divergence while releasing differentially private GMM models.
@@ I-A Our Contributions
    DP Mechanism Design via Privacy-Constrained KL-Divergence Minimization: Based on our analysis of model utility and privacy, we formulate an optimization problem that allocates privacy budgets and optimizes noise distributions by minimizing the expected KL divergence subject to a given $(\epsilon,\delta)$-DP constraint. While the problem is non-convex, we propose a low-complexity alternating optimization solution to compute a local optimum efficiently.
@@ I-B Organization and Notations
The remainder of this paper is organized as follows. We introduce the system model for differentially private parameter release of GMMs in Section II. In Section LABEL:sec3, we analyze the achievable DP and formulate the DP mechanism design problem. In Section LABEL:sec4, we introduce the proposed solution to the DP mechanism design problem. In Section LABEL:sec5, we present experimental results to evaluate the proposed method. Finally, this paper concludes in Section LABEL:sec6.
@@ II Differentially Private Distribution Release for GMMs
After estimating the parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$, we apply a post-processing step to satisfy given DP constraints before making these parameters public. The following definition of DP, introduced in [dwork2006calibrating, dwork2014algorithmic], is standard in the literature.
@@
Consider a randomized mechanism $M$ that takes a dataset $\mathcal{D}$ as input and outputs a query answer $M(\mathcal{D})$. Let $\epsilon>0$ and $\delta\in[0,1]$ be given parameters. $M$ is said to satisfy $(\epsilon,\delta)$-DP if, for any two adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ where $\mathcal{D}^{\prime}$ differs from $\mathcal{D}$ by altering the class label of exactly one data point [^2] For simplicity, we define DP in terms of an adjacent dataset $\mathcal{D}^{\prime}$ that alters exactly one class label from $\mathcal{D}$. The same principle extends to the variant where $\mathcal{D}^{\prime}$ replaces a single data point in $\mathcal{D}$ with a bounded alternative or removes/adds one data point from $\mathcal{D}$., the following inequality holds for all measurable sets $\mathcal{H}$:
@@
Toward this end, we assign each parameter a specific noise distribution. For the continuous-valued mean and covariance of each Gaussian component, we adopt a multivariate Gaussian mechanism and a Wishart mechanism [Jiang_Xie_Zhang_2016]. The perturbed means and covariances are given by
@@
where ${\bf w}_{k}$ and ${\bf W}_{k}$ are the noise terms added to the mean and covariance of the $k$-th component, respectively, $\boldsymbol{\Gamma}_{k}$ is the precision matrix of the Gaussian noise, $\mathcal{W}_{d}(\gamma_{k}{\bf I}_{d},d+1)$ is a Wishart distribution over $d\times d$ matrices with $d+1$ degrees of freedom and scale matrix $\gamma_{k}^{-1}{\bf I}_{d}$, and the parameter $\gamma_{k}>0$ controls the variance of the Wishart noise. Because Wishart random matrices are always positive definite, the construction in (6) ensures that the perturbed covariance remains feasible.

### v2

@@ I Introduction
In recent years, the remarkable success of data-driven artificial intelligence (AI) has spurred an increasing demand for the sharing and analysis of large-scale, multi-class, and high-dimensional datasets across a variety of domains, such as healthcare records, consumer transactions, and mobility traces. Organizations have recognized the potential of sharing data statistics to enhance data mining, improve public services, optimize recommendations, and facilitate data simulation [1]. However, sharing raw data or even their statistics raise significant privacy concerns, especially when sensitive attributes of individuals might be inferred. This tension underscores the need for privacy-preserving mechanisms that allow the release of data statistics without compromising personal or proprietary information [2].
@@
Among the various statistical modeling approaches, Gaussian Mixture Models (GMMs) stand out as a versatile tool for representing complex, multivariant, and multi-modal data distributions [3]. GMMs are widely applied in data mining, machine learning (ML), and pattern recognition. By modeling the overall distribution as a mixture of several Gaussian components, GMMs naturally capture latent subgroups or clusters in the data, such as different risk profiles in healthcare or distinct spending behaviors in retail. Furthermore, GMMs provide a compact and interpretable parameterization, including component means, covariances, and mixture weights (also known as categorical frequencies), which can be shared more efficiently than raw data records. As a motivating real-world application, GMMs are accurate in representing the statistics of energy consumption data. In particular, the logarithmic values of load demand profiles in Advanced Metering Infrastructure (AMI) data have been empirically shown to fit well to GMMs [4]. In the context of energy usage, these load demand profiles represent the amount of electricity consumed over time by a given user or household. When fitted to a GMM, the data naturally forms clusters, with each Gaussian component capturing different usage patterns (e.g., high, medium, and low usage times).
@@
Despite their advantages, releasing GMM parameters directly still poses privacy risks, particularly when the underlying dataset contains sensitive information. An adversary could potentially exploit small mixture components or extreme values to infer specific individuals’ records or their associated class labels. For instance, an adversary could inspect the component means of the GMM model fitted to the AMI dataset and detect changes that could identify the presence of a specific customer. Small perturbations in the model, such as flipping the label of a single customer, could significantly alter the estimated parameters, making it easier for the adversary to infer private information. To mitigate these risks, differential privacy (DP) [2, 5] offers a robust framework for quantifying and controlling the privacy loss resulting from the inclusion or exclusion of any individual record in a dataset. By introducing carefully calibrated noise, we can estimate mixture model parameters that satisfy strong privacy guarantees while still enabling meaningful statistical analysis.
@@
A central challenge in applying DP mechanisms to GMM fitting is maintaining fidelity between the original non-private data distribution and the released one under a DP constraint. Conventional DP mechanisms typically add artificial noise to perturb the model parameters in order to protect data privacy. However, this perturbation inevitably compromises the accuracy of the released GMM. For instance, the noise added to the Gaussian means could distort the true cluster centers, leading to incorrect clustering and a GMM that fails to represent the underlying data structure. Existing work [6] addressed differentially private GMMs by adding white Laplace noise to the mixture weights, means, and covariances individually. This approach models the accuracy of differentially private GMMs using the mean-squared errors (MSEs) of individual parameters.
@@
Rather than the MSEs that measure the accuracy of releasing individual parameters, we argue that distribution divergence, specifically Kullback–Leibler (KL) divergence, is another comprehensive metric for quantifying the fidelity of GMMs. KL divergence is a well-established statistical distance metric in information theory that quantifies how much one probability distribution diverges from another [7]. When used to characterize the utility of distribution release, KL divergence measures the “information” lost when substituting the true distribution with an approximate, privacy-preserving model. It is well-known that computing maximum likelihood estimation (MLE) is asymptotically equivalent to minimizing the KL divergence between the empirical and true distributions [8]. Thus, publishing a data distribution with minimal KL divergence ensures that the released model is as accurate as possible in the maximum-likelihood sense. Moreover, when applied to data simulation and augmentation for ML, KL divergence is closely related to the generalization ability of ML models. Extensive research has established PAC-Bayes bounds for quantifying the generalization error of ML models [9], which depends critically on the KL divergence between the training and testing data distributions. In this context, KL divergence serves as a powerful metric for evaluating how well an ML model trained on a given data distribution (e.g., a released GMM under privacy constraints) generalizes to the true data, thus providing an important tool for assessing the performance of data synthesis and augmentation in ML training.
@@
KL divergence provides a unified framework for characterizing the combined impact of noise perturbation on the accuracy of the released GMM, taking into account the effects of noise on all model parameters. However, characterizing the KL divergence between the original and DP-preserved distributions, particularly when applying DP mechanisms to the released GMM parameters, remains an open problem. In this work, we show that the KL divergence for GMM parameter release exhibits a closed-form expression under specific DP mechanisms, thereby opening new research avenues for designing DP mechanisms that minimize KL divergence while releasing differentially private GMM models.
@@ I-A Our Contributions
    DP Mechanism Design via Privacy-Constrained KL-divergence Minimization: Based on our analysis of model utility and privacy, we formulate an optimization problem that allocates privacy budgets and optimizes noise distributions by minimizing the expected KL divergence subject to a given $(\epsilon,\delta)$-DP constraint. Crucially, this formulation makes the privacy–utility trade-off transparent: the privacy budget is reallocated across DP randomization parameters during the alternating updates, directly exposing how noise levels and budget shares affect the utility. While the problem is non-convex, we propose a low-complexity alternating optimization solution to compute a local optimum efficiently.
@@ I-B Organization and Notations
The remainder of this paper is organized as follows. We introduce the system model for differentially private parameter release of GMMs in Section II. In Section III, we analyze the achievable DP and formulate the DP mechanism design problem. In Section IV, we introduce the proposed solution to the DP mechanism design problem. In Section V, we present experimental results to evaluate the proposed method. Finally, this paper concludes in Section VI.
@@ II Differentially Private Distribution Release for GMMs
After estimating the parameters $\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$, we apply a post-processing step to satisfy given DP constraints before making these parameters public. The following definition of DP, introduced in [2, 5], is standard in the literature.
@@
Consider a randomized mechanism $M$ that takes a dataset $\mathcal{D}$ as input and outputs a query answer $M(\mathcal{D})$. Let $\epsilon>0$ and $\delta\in[0,1]$ be given parameters. $M$ is said to satisfy $(\epsilon,\delta)$-DP if, for any two adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ where $\mathcal{D}^{\prime}$ differs from $\mathcal{D}$ by altering the class label of exactly one data point, the following inequality holds for all measurable sets $\mathcal{H}$:
@@
Remark 1.
Definition 1 adopts label-level DP, where an adjacent dataset $\mathcal{D}^{\prime}$ alters exactly one class label. Record-level DP is also possible under bounded features, with adjacency defined by adding, removing, or replacing one data point. Section IV-C shows how our approach adapts to record-level DP with the same design principles.
@@
Toward this end, we assign each parameter a specific noise distribution. For the continuous-valued mean and covariance of each Gaussian component, we adopt a multivariate Gaussian mechanism and a Wishart mechanism [10]. The perturbed means and covariances are given by
@@
where ${\bf w}_{k}$ and ${\bf W}_{k}$ are the noise terms added to the mean and covariance of the $k$-th component, respectively, $\boldsymbol{\Gamma}_{k}$ is the precision matrix of the Gaussian noise, $\mathcal{W}_{d}(\gamma_{k}^{-1}{\bf I}_{d},d+1)$ is a Wishart distribution over $d\times d$ matrices with $d+1$ degrees of freedom and scale matrix $\gamma_{k}^{-1}{\bf I}_{d}$, and the parameter $\gamma_{k}>0$ controls the variance of the Wishart noise. Because Wishart random matrices are always positive definite, the construction in (6) ensures that the perturbed covariance remains feasible.